OpenStudy (anonymous):

Multiply: z squared plus 6 times z plus 5 all over 3. • 9 times x all over the quantity z squared plus 5 times z, end quantity.

OpenStudy (anonymous):

Can someone please help? There's a few more like it I don't understand..

OpenStudy (anonymous):

\[\frac{z^2+6z+5}{3}*\frac{9x}{z^2+5z}\]

OpenStudy (anonymous):

do you know how to fator Pedro?

OpenStudy (anonymous):

factor*

OpenStudy (anonymous):

Somewhat. I use online school and it doesn't teach very well, but I'll give it a shot.

OpenStudy (anonymous):

try factoring the top of the first and the bottom of the second

OpenStudy (anonymous):

would it be (5 + z)(1 + z)?

OpenStudy (anonymous):

For the top?

OpenStudy (anonymous):

yes and the botom of the second has just a GCF

OpenStudy (anonymous):

So it would stay 3?

OpenStudy (anonymous):

just leave the bottom as 3. you just want to expand the problem first

OpenStudy (anonymous):

Ok, makes sense so far :) Would you factor the second half next?

OpenStudy (anonymous):

so for \[z^2+5z\] what is your Greatest common factor, or what can you pull out

OpenStudy (anonymous):

the 5?

OpenStudy (anonymous):

not quite, what does \[z^2\] and \[5z\] have in common

OpenStudy (anonymous):

the z

OpenStudy (anonymous):

so you would combing the two?

OpenStudy (anonymous):

yes so the gcf is z, if you pull z out of each what are you left with z(? + ?)

OpenStudy (anonymous):

um, 3 and 2?

OpenStudy (anonymous):

so greatest common factors work sorta like factoring only you don't have a 3terms so like \[x^2+15x\] you can see there is an x in both of them so if you pul out an x, what are you left with well \[x*x=x^2\] and\[15*x=15x\] so if you get rid of one x in both of these and bring it out front you end up with \[x(x+15)\] if you were to multiply it back out you'd get what you originally had

OpenStudy (anonymous):

so for yours \[z*z=z^2\] \[5*z=5z\] if you do the same that i did above what do you get

OpenStudy (anonymous):

z(x^2+5z) ?

OpenStudy (anonymous):

z^2 sorry

OpenStudy (anonymous):

if you multiplied that back out would you get the same thing you had before?

OpenStudy (anonymous):

you forgot to get rid of the z you pulled out of the original equation =o

OpenStudy (anonymous):

so basically to check if you did it correctly use the distibutive property \[a(b+c)=ab+ac\] \[z(z^2+5z)=z^3+5z^2\] o no! thats not what you hadbefore!

OpenStudy (anonymous):

\[z*z=z^2\] \[5*z=5z\] if you pull a z out of each \[z(z+5)=z^2+5\]

OpenStudy (anonymous):

5z*

OpenStudy (anonymous):

so now you have \[\frac{(z+1)(z+5)}{3}*\frac{9x}{z(z+5)}\]

OpenStudy (anonymous):

when multiplying what can you do with like terms in the numerator and denominator?

OpenStudy (anonymous):

Oh, so then you would cancel the like terms out!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

z+5 can cancel out, right?

OpenStudy (anonymous):

\[\frac{z+1}{3}*\frac{9x}{z}\] then multiply them

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you'll get \[\frac{(z+1)9x}{3z}\]

OpenStudy (anonymous):

would there be restrictions?

OpenStudy (anonymous):

like z can't equal 0 or -5?

OpenStudy (anonymous):

umm the denominator cannot equal zero so the only time that it would equal zero is if z itself was zero

OpenStudy (anonymous):

ok and there isn't more simplifying? Like the 3z and z+1?